OpenStudy (moonlitfate):

In a completely elastic collision, m1 = 10 kg is traveling from left to right at a speed 10 m/s and collides with m2 = 5 kg traveling at a speed of 14 m/s traveling toward it. After collision, what is the velocity of the first ball to the nearest tenth of a m/s, if the second ball's velocity is 14 m/s? I ended up getting 10.0 m/s, but it was wrong. Can someone show me where I went wrong?

OpenStudy (moonlitfate):

@zephyr141 -- would it help if I showed you the steps that I took to get my answer first?

OpenStudy (zephyr141):

no i see what you did

OpenStudy (zephyr141):

you set m1's velocity to negative.

OpenStudy (zephyr141):

m1 is traveling from left TO RIGHT so it should be positive

OpenStudy (zephyr141):

\[(10*10)+[5*(-14)]=10v_{1}+(5*14)\] i'm assuming here that the second 14 m/s is the final velocity of ball 2

OpenStudy (moonlitfate):

This is what I wrote down for the information given (just trying to understand this so sinks in): \[m _{1} = 10 kg \]\[m_{2} = 5 kg\] \[v_{1} = 10 m/s\]\[v_{2}= -14 m/s\]\[v_{2}^{f} = -14 m/s\]

OpenStudy (zephyr141):

is the that all written down word for word? if so then i think that the final velocity of ball 2 is positive 14m/s.

OpenStudy (zephyr141):

i mean the question

OpenStudy (moonlitfate):

Yes. That is exactly how it was written.

OpenStudy (zephyr141):

then i think the final velocity, or v2 final, is going to be positive 14 m/s

OpenStudy (moonlitfate):

Oh... wait, would that because because after the collision.. the 2nd ball would go in the opposite direction?

OpenStudy (zephyr141):

yeah. you typed it out exactly as it was given, as you said, so i think it's going in the positive direction after the collision.